Which is the most polar bond in the following list? OS-O OP-H OOH OC-s O1
Answer #1

Answer;;;;; O-H is most polar bond than other there

Let's the electronegativity value of the given elements in pauling scale are,

C= 2.55

N=3.04

P=2.19

S=2.58

O=3.44

F=3.98

I = 2.66

H=2.2

we calculate the  electronegativity value because which is most different then other ,this is most polar

S-O =3.44 -2.58=.86

P-H =2.2 -2.19=0.01

O-H= 3.44 -2.2=1.24

C-S =2.55 - 2.58=0..03

I-I = 2.66 -2.66=0

we notice there O- H is highest electronegativity different so O-H is most polar bond than other element.